# Elk: Exploring the Efficiency of Inter-core Connected AI Chips with Deep Learning Compiler Techniques

### v1

@@ 4.4. Preload Order Permutation
First, reordering helps mitigate interconnect contention. As the interconnect traffic pressure fluctuates (see §$\S$3.3), the reordering opportunistically reschedules heavy preload traffic to avoid “rush hours” on the interconnect.
@@
Generate valid preload orders. Elk enumerates all valid preload orders by scanning through all operators following the inductive operator scheduling order (§$\S$4.2) and incrementally picking the next operator to preload in each step.
@@
For each generated preload order, Elk invokes the operator scheduling pass in §$\S$4.2, forming a $O(C^{H}PK^{2}N)$ search space. Elk picks the best end-to-end plan among all preload orders.
@@ 4.5. Mapping to Hardware
[…]
@@ 5. Implementation details
For each operator, we enumerate all possible partition plans by representing each plan as a list of integers. For instance, ¡90,9¿ evenly slices each dimension of a 2-dimension operator into 90 and 9 parts, forming 90$\times$9=810 tiles. For each plan, Elk decides the mapping of each tile to each core. It uses different mapping strategies for different network topologies. Elk currently targets ICCA chips with two popular network topologies: all-to-all network and mesh network. For chip with all-to-all network, Elk sequentially maps all tiles, as core locations do not impact the inter-core data transfer cost. For chip with $N$-dimensional mesh network, Elk chooses from plans that partition an operator along at most $N$ dimensions, so it can map each partitioned dimension to a mesh dimension. Then, Elk uses dimension-order routing (Jouppi et al., 2023; He et al., 2025) to maximize the all-reduce bandwidth. Besides the two topologies that are used by most ICCA chips today, Elk is scalable to support other topologies.
@@
Figure 16. Elk compile time for varied model/batch sizes.
Simulation framework. To conduct sensitivity analysis and design space exploration, we build an event-driven simulator for ICCA chips, which simulates all cores, network links, and off-chip HBM accesses. For each core, we simulate a local SRAM, a compute pipeline, and a network agent that sends/receives data to/from other cores. For each network link, we model its latency and bandwidth (Abdelhafez et al., 2019; Alexandrov et al., 1995). Based on the execution plan generated by Elk, we derive the simulation events at tile granularity, including computing a tile on a core, transferring a tile over a specific network link, and fetching a tile from the off-chip HBM. Each core/link maintains its event queue to execute its events sequentially. For an all-to-all network, we model HBM controllers as dedicated nodes in the network (see §$\S$2.1). For a mesh network, we attach HBM controllers to the edges of the mesh grid. To simulate a multi-chip system, we track the in-flight inter-chip transfer events and cap their total bandwidth. We also use our real IPU-based emulator to validate our simulator.
@@ 6.1. Experimental Setup
Baselines. As there is no open-sourced compiler for ICCA chip with HBM, we conduct an ablation study by creating two baselines that extend SOTA compilers for ICCA chips (Liu et al., 2024) to support HBM, and an Elk variant that disables preload reordering (§$\S$4.4). We also compare Elk to an ideal roofline. In brief, we compare these designs:
@@ 6.2. End-to-end Performance
Figure 17 shows the per-token generation latency of LLM decoding on our emulator. On average, Elk-Full outperforms Basic by 1.87$\times$ (up to 1.93$\times$), Static by 1.37$\times$ (up to 1.49$\times$), and achieves 94.84% of the ideal performance. The performance of Elk also scales well with increasing batch size and sequence length. Notably, Gemma2-27B and Llama2-70B can achieve latencies similar to those of smaller LLMs, since they use Grouped-Query Attention (Ainslie et al., 2023).
Inference latency breakdown. In Figure 18 (a), we break total time into four categories: 1 preload (HBM is loading), 2 execute (cores are computing/sending data), 3 overlapped preload & execute, and 4 interconnect (HBM/cores are stalled by interconnect contention). We only show batch size 32 and sequence length 2048 due to space limits. Basic always poorly overlaps preload and per-core execution. By preloading more operators, Static increases the overlap time by 11.26$\times$, but is limited by fixed preload and execution space sizes. Elk-Dyn overlaps better by adjusting the on-chip memory allocation based on operators’ demands, but suffers from interconnect congestion and misses preload opportunities (when the available preload space is too small for the next operator, but can fit a future operator). By reordering preloads with an average edit distance of 2.9 steps, Elk-Full eliminates 87.65% of interconnect congestion overhead over Elk-Dyn. Elk-Full also reduces the non-overlapped preload time to 0.037% of the total, because of reduced on-chip memory contention.
@@ 6.4. Design Space Exploration for ICCA Chips
(3) Elk enables scalable performance for ML inference workloads as we scale the ICCA chip. In Figure 23, we change the number of cores while setting the HBM bandwidth to 2.7GBps/core to match prior setups. Elk-Full significantly outperforms other designs regardless of core counts. Elk-Full reduces the average latency by 1.71$\times$ over Basic and 1.36$\times$ over Static. We also examine DiT-XL, a state-of-the-art stable diffusion model, on one ICCA chip (up to 1472 cores). Elk-Full’s benefit on DiT-XL is less obvious than on LLMs, since DiT-XL is compute-intensive and less affected by preload efficiency. However, Elk-Full still outperforms other designs on DiT-XL and achieves near-ideal performance.
(4) ICCA chips can also benefit ML training by properly tuning the compute, communication, and off-chip memory access. In Figure 24, we examine the forward pass of training Llama2-13B with varied available FLOPS and interconnect/HBM bandwidths (the backward pass has similar trends). Unlike decoding, training is compute-intensive, scaling only interconnect/HBM bandwidth has little impact. With 400GB/s HBM bandwidth, it is sufficient to fulfill more than 650 TFLOPS. Thus, for compute-intensive workloads, the ICCA chips should focus on scaling the FLOPS, and can therefore be paired with cheaper memory (e.g., GDDR/LPDDR/DDR) to reduce manufacturing costs. Note that the achieved FLOPS is often lower than the peak FLOPS of the hardware, because only MatMul operators with perfect shapes can fully utilize the FLOPS of specialized tensor cores.
@@ 7. Discussion and Future Work
Apply Elk to GPUs. The latest NVIDIA GPU also uses inter-core links to connect its stream multiprocessors (SMs) (hop, 2022). It groups SMs into clusters. SMs in the same cluster are connected via direct inter-SM links, while different clusters can only exchange data via the global L2 cache. On current GPUs like H100 (NVIDIA H100 Tensor Core GPU, [n. d.]), the aggregated inter-SM bandwidth is close to the HBM bandwidth, so it will suffer from significant interconnect contention ( 2 in Figure 2). As future work, we wish to extend Elk to GPU and investigate the design space for optimizing GPU’s interconnect architecture.
@@ 8. Related Work
ML optimizations with operator fusion. DL frameworks (Google, 2023; Zheng et al., 2022c; Shi et al., 2023; Cai et al., 2025; Dao et al., 2022) improve on-chip data reuse by fusing multiple operators. For example, SoMa (Cai et al., 2025) explores fusion opportunities to reuse intermediate tensors between operators in a global on-chip buffer with limited size. As ICCA chips have large distributed SRAM (e.g., up to 900MB per chip) that can buffer an entire intermediate tensor, they can reuse it between operators without fusion. However, ICCA chips face unique challenges in distributed SRAM allocation, inter-core interconnect contention, and their impacts on HBM preload. Thus, Elk focuses on the interplay of these performance challenges. For ICCA chips with less SRAM, Elk can still support fusion by treating each fused operator as one operator. For example, we can fuse two consecutive MatMul operators into one operator (Dao et al., 2022), which is treated as one operator with three input tensors in Elk.
@@
Compilers for ICCA chips. Dataflow compilers optimize DL execution on interconnected cores by mapping operators to a pipeline (Prabhakar et al., 2017; Zheng et al., 2022a; Cai et al., 2023; Lu et al., 2017; Zhang et al., 2021, 2019). While they optimize on-chip execution and communication, they use off-chip HBM differently. SambaNova SN40L (Prabhakar et al., 2024) swaps parameters between HBM and DDR to serve different expert models. Tenstorrent (Bahnas, 2024) optimizes intra-operator tile size to reduce the total off-chip memory access volume. T10 (Liu et al., 2024) optimizes on-chip execution without considering off-chip memory, so it cannot support LLMs that exceed on-chip capacity (e.g., $<$1GB per chip). Specifically, T10 can only optimize the placement of a fixed set of operators in a fixed amount of SRAM. However, as an ICCA chip loads new operators from the HBM and frees old ones from the SRAM, the set of operators in the SRAM changes dynamically. In addition, the HBM accesses will interfere with the on-chip execution due to the SRAM and NoC contentions. As T10 cannot consider all performance tradeoffs together, even after we extend T10 to support HBM, the extended design (i.e., Static in §6.1) still performs poorly on ICCA chips with HBM. To maximize performance for generic ICCA chips, Elk considers the interplay among off-chip HBM, on-chip SRAM, inter-core interconnect, and per-core execution on top of the on-chip execution optimizations enabled by existing compilers for ICCA chips.
@@
Distributed model execution. To run DL models on distributed nodes, prior works overlap computation with inter-device communication. They optimize inter-chip collective communication (Chen et al., 2024; Rashidi et al., 2022; Jangda et al., 2022; Wang et al., 2022), device grouping on various network topologies (Zheng et al., 2022b; Um et al., 2024; Li et al., 2022; Miao et al., 2022), and workload collocation in device clusters (Xue et al., 2023, 2024; Feng et al., 2025). Elk targets intra-chip optimization and faces other challenges. Besides compute and communication, Elk also considers HBM data accesses and their impacts on the usages of per-core SRAM and inter-core links. Elk is compatible with various distributed execution frameworks and different parallelism types when using multiple ICCA chips, where Elk optimizes the execution of each ICCA chip.
@@
ML- and SAT-based compiler optimizations. Existing studies like Autocomp (Hong et al., 2024) use LLMs to optimize kernel code for tensor accelerators, demonstrating significant performance advantages over vendor libraries. The code generation in Elk (§5) can leverage these works to further optimize the per-core tile computation, by replacing the code templates from vendor libraries with the kernels optimized by LLMs. Elk’s scheduling algorithm (§4.2–§4.4) does not rely on LLMs, since it can already find execution plans with near-Ideal performance (see §6.2) in short compile time (see Figure 16) using the host CPU. Both Elk’s comprehensive optimization space and its pruning techniques based on hardware information (e.g., SRAM size, core count) contribute to its success. Other prior works (Vanhoucke and Coelho, 2016; Hebrard, 2012) use SAT solvers to solve scheduling problems. However, it is inefficient to use SAT solvers in Elk, because their runtime grows exponentially with the number of boolean variables to solve (Gong and Zhou, 2017). For example, to solve Elk’s allocation problem in §4.3 using SAT, we need to assign one boolean variable to each possible execution plan of each operator to indicate whether this plan is selected. Thus, the solver’s runtime grows exponentially with the total number of possible plans of all operators. In comparison, the complexity of our allocation algorithm in §4.3 grows linearly with the total number of plans (i.e., $O(PK)$ for $K$ operators that each has $P$ possible plans).
@@ 9. Conclusion
We study the performance trade-offs of generic ICCA chips that support off-chip memory, and develop a DL compiler framework Elk to explore the efficiency of ICCA chips. Elk also enables design space exploration of ICCA chip architecture. We demonstrate the capability of Elk using both an ICCA emulator and simulator.
Appendix A Proofs of Lemma 4.1 and Theorem 4.2
Lemma 4.1 (Base case).
Proof.
Theorem 4.2 (Inductive step).
[…]
Proof.
[…]
Then, we prove the existence of $T^{i}_{\textit{{s}-{exe}}}$ that minimizes $T_{end}-T^{i}_{\textit{{s}-{exe}}}$. Suppose by contradiction that $T_{end}-T^{i+1}_{\textit{{s}-{exe}}}$ is minimized but there is no $T^{i}_{\textit{{s}-{exe}}}$ that minimizes $T_{end}-T^{i}_{\textit{{s}-{exe}}}$. Since our inductive step explored all $T^{i}_{\textit{{s}-{exe}}}$ values by enumerating all possible preload numbers, the only possible case is that we must explore more preload numbers to find the global $\max(T^{i}_{\textit{{s}-{exe}}})$, or the global $\max(T^{i}_{\textit{{e}-{exe}}})$ is greater than $T^{i+1}_{\textit{{s}-{exe}}}$, which means $T^{i+1}_{\textit{{s}-{exe}}}$ can be larger. This is a contradiction since $T_{end}-T^{i+1}_{\textit{{s}-{exe}}}$ is already minimized, e.g., $T^{i+1}_{\textit{{s}-{exe}}}$ is already maximized. ∎

### v2

@@ 4.4. Preload Order Permutation
First, reordering helps mitigate interconnect contention. As the interconnect traffic pressure fluctuates (see $\S$3.3), the reordering opportunistically reschedules heavy preload traffic to avoid “rush hours” on the interconnect.
@@
Generate valid preload orders. Elk enumerates all valid preload orders by scanning through all operators following the inductive operator scheduling order ($\S$4.2) and incrementally picking the next operator to preload in each step.
@@
For each generated preload order, Elk invokes the operator scheduling pass in $\S$4.2, forming a $O(C^{H}PK^{2}N)$ search space. Elk picks the best end-to-end plan among all preload orders.
@@ 4.5. Mapping to Hardware
Figure 15. The abstracted device programming model of Elk.
The execution plan generated by Elk specifies all operator’s preload order and each operator’s partition plans. Elk maps the plan to an abstracted programming model, which can be applied to generic ICCA chips with off-chip memory. As shown in Figure 15, Elk abstracts two key device functions that are generated during compilation. (1) preload_async(op=i) commands all cores to request Opii’s data from HBM based on the preload-state partition plan. (2) execute(op=i) runs Opii on all cores based on the execute-state plan.
@@ 5. Implementation details
[…]
@@
Figure 16. Elk compile time for varied model/batch sizes.
@@
Simulation framework. To conduct sensitivity analysis and design space exploration, we build an event-driven simulator for ICCA chips, which simulates all cores, network links, and off-chip HBM accesses. For each core, we simulate a local SRAM, a compute pipeline, and a network agent that sends/receives data to/from other cores. For each network link, we model its latency and bandwidth (Abdelhafez et al., 2019; Alexandrov et al., 1995). Based on the execution plan generated by Elk, we derive the simulation events at tile granularity, including computing a tile on a core, transferring a tile over a specific network link, and fetching a tile from the off-chip HBM. Each core/link maintains its event queue to execute its events sequentially. For an all-to-all network, we model HBM controllers as dedicated nodes in the network (see $\S$2.1). For a mesh network, we attach HBM controllers to the edges of the mesh grid. To simulate a multi-chip system, we track the in-flight inter-chip transfer events and cap their total bandwidth. We also use our real IPU-based emulator to validate our simulator.
@@ 6.1. Experimental Setup
Baselines. As there is no open-sourced compiler for ICCA chip with HBM, we conduct an ablation study by creating two baselines that extend SOTA compilers for ICCA chips (Liu et al., 2024) to support HBM, and an Elk variant that disables preload reordering ($\S$4.4). We also compare Elk to an ideal roofline. In brief, we compare these designs:
@@ 6.2. End-to-end Performance
[…]
Figure 19. Per-token latency at varied HBM bandwidths.
[…]
@@ 6.4. Design Space Exploration for ICCA Chips
(4) ICCA chips can also benefit ML training by properly tuning the compute, communication, and off-chip memory access. In Figure 24, we examine the forward pass of training Llama2-13B with varied available FLOPS and interconnect/HBM bandwidths (the backward pass has similar trends). Unlike decoding, training is compute-intensive, scaling only interconnect/HBM bandwidth has little impact. With 400GB/s HBM bandwidth, it is sufficient to fulfill more than 600 TFLOPS. Thus, for compute-intensive workloads, the ICCA chips should focus on scaling the FLOPS, and can therefore be paired with cheaper memory (e.g., GDDR/LPDDR/DDR) to reduce manufacturing costs. Note that the achieved FLOPS is often lower than the peak FLOPS of the hardware, because only MatMul operators with perfect shapes can fully utilize the FLOPS of specialized tensor cores.
@@ 7. Discussion and Future Work
Apply Elk to GPUs. The latest NVIDIA GPU also uses inter-core links to connect its stream multiprocessors (SMs) (hop, 2022). It groups SMs into clusters. SMs in the same cluster are connected via direct inter-SM links, while different clusters can only exchange data via the global L2 cache. On current GPUs like H100 (NVIDIA H100 Tensor Core GPU, 2024), the aggregated inter-SM bandwidth is close to the HBM bandwidth, so it will suffer from significant interconnect contention ( in Figure 2). As future work, we wish to extend Elk to GPU and investigate the design space for optimizing GPU’s interconnect architecture.
@@ 8. Related Work
ML optimizations with operator fusion. DL frameworks (Google, 2023; Zheng et al., 2022c; Shi et al., 2023; Cai et al., 2025; Dao et al., 2022; Luo et al., 2025) improve on-chip data reuse by fusing multiple operators. For example, SoMa (Cai et al., 2025) explores fusion opportunities to reuse intermediate tensors between operators in a global on-chip buffer with limited size. As ICCA chips have large distributed SRAM (e.g., up to 900MB per chip) that can buffer an entire intermediate tensor, they can reuse it between operators without fusion. However, ICCA chips face unique challenges in distributed SRAM allocation, inter-core interconnect contention, and their impacts on HBM preload. Thus, Elk focuses on the interplay of these performance challenges. For ICCA chips with less SRAM, Elk can still support fusion by treating each fused operator as one operator. For example, we can fuse two consecutive MatMul operators into one operator (Dao et al., 2022), which is treated as one operator with three input tensors in Elk.
@@
Compilers for ICCA chips. Dataflow compilers optimize DL execution on interconnected cores by mapping operators to a pipeline (Prabhakar et al., 2017; Zheng et al., 2022a; Cai et al., 2023; Lu et al., 2017; Zhang et al., 2021, 2019; He et al., 2025). While they optimize on-chip execution and communication, they use off-chip HBM differently. SambaNova SN40L (Prabhakar et al., 2024) swaps parameters between HBM and DDR to serve different expert models. Tenstorrent (Bahnas, 2024) optimizes intra-operator tile size to reduce the total off-chip memory access volume. T10 (Liu et al., 2024) optimizes on-chip execution without considering off-chip memory, so it cannot support LLMs that exceed on-chip capacity (e.g., <<1GB per chip). Specifically, T10 can only optimize the placement of a fixed set of operators in a fixed amount of SRAM. However, as an ICCA chip loads new operators from the HBM and frees old ones from the SRAM, the set of operators in the SRAM changes dynamically. In addition, the HBM accesses will interfere with the on-chip execution due to the SRAM and NoC contentions. As T10 cannot consider all performance tradeoffs together, even after we extend T10 to support HBM, the extended design (i.e., Static in §6.1) still performs poorly on ICCA chips with HBM. To maximize performance for generic ICCA chips, Elk considers the interplay among off-chip HBM, on-chip SRAM, inter-core interconnect, and per-core execution on top of the on-chip execution optimizations enabled by existing compilers for ICCA chips.
@@
Distributed model execution. To run DL models on distributed nodes, prior works overlap computation with inter-device communication. They optimize inter-chip collective communication (Chen et al., 2024; Rashidi et al., 2022; Jangda et al., 2022; Wang et al., 2022), device grouping on various network topologies (Zheng et al., 2022b; Um et al., 2024; Li et al., 2022; Miao et al., 2022), and workload collocation in device clusters (Xue et al., 2023b, 2024; Feng et al., 2025; Xue et al., 2023a). Elk targets intra-chip optimization and faces other challenges. Besides compute and communication, Elk also considers HBM data accesses and their impacts on the usages of per-core SRAM and inter-core links. Elk is compatible with various distributed execution frameworks and different parallelism types when using multiple ICCA chips, where Elk optimizes the execution of each ICCA chip.
@@
ML- and SAT-based compiler optimizations. Existing studies like Autocomp (Hong et al., 2024) use LLMs to optimize kernel code for tensor accelerators, demonstrating significant performance advantages over vendor libraries. The code generation in Elk5) can leverage these works to further optimize the per-core tile computation, by replacing the code templates from vendor libraries with the kernels optimized by LLMs. Elk’s scheduling algorithm (§4.2–§4.4) does not rely on LLMs, since it can already find execution plans with near-Ideal performance (see §6.2) in short compile time (see Figure 16) using the host CPU. Both Elk’s comprehensive optimization space and its pruning techniques based on hardware information (e.g., SRAM size, core count) contribute to its success. Other prior works (Vanhoucke and Coelho, 2016; Hebrard, 2012) use SAT solvers to solve scheduling problems. However, it is inefficient to use SAT solvers in Elk, because their runtime grows exponentially with the number of boolean variables to solve (Gong and Zhou, 2017). For example, to solve Elk’s allocation problem in §4.3 using SAT, we need to assign one boolean variable to each possible execution plan of each operator to indicate whether this plan is selected. Thus, the solver’s runtime grows exponentially with the total number of possible plans of all operators. In comparison, the complexity of our allocation algorithm in §4.3 grows linearly with the total number of plans (i.e., $O(PK)$ for $K$ operators that each has $P$ possible plans).
@@ 9. Conclusion
We study the performance trade-offs of generic ICCA chips that support off-chip memory, and develop a DL compiler framework Elk to explore the efficiency of ICCA chips. Elk also enables design space exploration of ICCA chip architecture. We demonstrate the capability of Elk using both an ICCA emulator and a simulator.
Acknowledgements.
We thank the anonymous reviewers at MICRO’25 for their insightful feedback. We thank Michael Wang and Benjamin Reidys from the Systems Platform Research Group (Illinois PlatformX) at UIUC for proofreading our paper. This work was partially supported by the Hybrid Cloud and AI program at the IBM-Illinois Discovery Accelerator Institute (IIDAI), and NSF under the grants CAREER CNS-2144796, CCF-2107470, and CCF-1919044.
Appendix A Artifact Appendix
A.1. Abstract
In this artifact, we provide the source code of Elk’s compilation, simulation, and evaluation framework. Then, we guide readers to explore how Elk improves the model serving performance on a variety of ICCA chips (i.e., Figure 17-24 in this paper). To run this artifact, please use a Linux machine with at least 200 GB of main memory and at least 20 GB of disk space.
A.2. Artifact Checklist (Meta-Information)
  • Algorithm: Inductive tensor operator scheduling, cost-aware on-chip memory allocation, and ICCA chip design space exploration.
  • Neural Network Models: Llama2-13B, Gemma2-27B, OPT-30B, Llama2-70B, and DiT-XL. Their execution graphs are included in the repo.
  • Run-time environment: Ubuntu 20.04 or newer, Python 3.10.
  • Metrics: Execution time, hardware utilization.
  • Output: Trace files and result figures.
  • Experiments: Generate experiments using supplied scripts.
  • How much main memory required (approximately): 200 GB
  • How much disk space required (approximately): 20 GB
  • How much time to prepare workflow (approximately): 10 minutes
  • How much time to complete experiments (approximately): 30 hours on a machine with 64 CPU threads and 200 GB main memory.
  • Publicly available: Yes
  • Archived (provide DOI): 10.5281/zenodo.16541972
A.3. Description
A.3.1. How to Access
The source code can be downloaded from Zenodo at https://doi.org/10.5281/zenodo.16541972. For the latest version, you can access our GitHub repository: https://github.com/platformxlab/elk.git.
A.3.2. Hardware Dependencies
The Elk simulation and evaluation framework can run on any x86 machine with at least 200 GB of main memory and at least 20 GB of disk space.
A.3.3. Software Dependencies
The framework needs a Linux environment (preferably Ubuntu) with Python 3.10 installed.
A.4. Installation
frame=tb, language=bash, aboveskip=3mm, belowskip=3mm, showstringspaces=false, columns=flexible, basicstyle=, numbers=none, numbersep=0pt, numberstyle=, keywordstyle=, commentstyle=, stringstyle=, breaklines=true, breakatwhitespace=true, tabsize=5, classoffset=0, keywordstyle=,
  1. (1)
    Start by downloading the Elk artifact from GitHub:
    git clone https://github.com/platformxlab/elk.git
    cd elk
  2. (2)
    Please make sure all prerequisites are successfully installed:
    sudo add-apt-repository ppa:deadsnakes/ppa
    sudo apt update
    sudo apt install python3.10 tmux -y
    curl -sS https://bootstrap.pypa.io/get-pip.py | python3.10
    python3.10 -m pip install -r requirements.txt
A.5. Experiment Workflow
To compile DL models into programs and obtain program execution traces from the Elk simulator, we provide a one-click script ”benchmark_scripts/generate_data_from_sim.py” for you to launch all test cases in one place. However, the script may take more than 30 hours to finish, so we recommend using tmux:
tmux
Then, run the one-click script within the new tmux terminal:
python3.10 benchmark_scripts/generate_data_from_sim.py
To return from the tmux terminal without pausing the script, press ”Ctrl+B” and then press ”D” on your keyboard. To attach back to the original tmux terminal where the script is running, use:
tmux attach -t 0
For more tips on using tmux, refer to https://tmuxcheatsheet.com.
A.5.1. Handle Errors
If the script encounters an error, the most common cause is that the artifact runs on too many CPU cores and overflows the main memory. In such events, (1) go to ”launch.py”, (2) change the ”CORE_REDUCE” macro in line 22 to a larger value (e.g., CORE_REDUCE=8), and (3) rerun the script:
python3.10 benchmark_scripts/generate_data_from_sim.py
The script should automatically skip any completed test cases and resume from the failed one.
A.6. Evaluation and Expected Results
After the completion of all experiments, please run the following script to evaluate the results:
./run_artifact_eval_graph_gen.sh
This script gathers all data from the execution trace and draws all figures. To verify the results, one can compare the generated figures with those in the paper.
A.7. Methodology
Submission, reviewing and badging methodology: